# ORCID: https://orcid.org/0000-0001-5105-1463

### v2

@@ Abstract
The ongoing controversy surrounding transformative agreements, which aim to transition journal publishing to full open access, highlight the need for large-scale studies assessing the impact of these agreements on hybrid open access. By combining publicly available data from various sources, including cOAlition S Journal Checker, Crossref, and OpenAlex, this study presents a novel approach that analyses over 700 agreements. Results suggest a strong growth in open access between 2018 and 2022 from 4.3% to 15%. During this period, 11,189 hybrid journals provided open access to 742,369 out of 8,146,958 articles, representing a five-year open access proportion of 9.1%. Authors who could make use of transformative agreements at the time of publication contributed 328,957 open access articles. In 2022, 143,615 out of 249,511 open access articles in hybrid journals (58%) were enabled by transformative agreements. This trend was largely driven by the three commercial publishers Elsevier, Springer Nature, and Wiley, but the open access uptake varied substantially across journals, publishers, disciplines, and country affiliations. In particular, the OECD and BRICS areas revealed different publication trends. In conclusion, this study suggests that current levels of implementation of transformative agreements is insufficient to bring about a large-scale transition to full open access.
@@ 1 Introduction
As a consequence, libraries and their consortia began to develop licensing strategies aimed at avoiding such ‘double dipping’ scenarios, in which well-established commercial publishers gain twice from reading and open access publishing (Mittermaier, 2015), as well as to increase publisher-provided immediate open access (Björk & Solomon, 2014; Schimmer et al., 2015). These considerations resulted in transformative agreements [^1] In this paper I use the term “transformative agreement”, addressing also offsetting, read-and-publish or publish-and-read deals, and other variants (Borrego et al., 2021; Hinchliffe, 2019). Although the term is critised as misleading and not useful to describe the different types of open access agreements between library consortia and commercial publishers (Babini et al., 2022), it is widely used in policy discussions and in the research literature., which cover a broad range of contracts between library consortia and publishers from the mid-2010s onwards, where institutional spending for subscriptions and open access publishing are considered together (Borrego et al., 2021; Hinchliffe, 2019). Transformative agreements seek to control costs while allowing a transitional phase for publishing more open access articles. Similar to big deals, transformative agreements mainly bundle hybrid and subscription-only journals from commercial publishers, but aim at a higher degree of transparency than previous big deals, where contracts including payments were confidential (Bergstrom et al., 2014).
@@
By the end of 2023, many transformative agreements had been implemented, but the interim outcomes were mixed. The ESAC Transformative Agreement Registry [^2] https://esac-initiative.org/about/transformative-agreements/agreement-registry/, the largest resource for library consortia to publicly share their agreements, recorded more than 800 transformative agreements. These agreements resulted in the publication of up to 900.000 open access articles published in both fully open access and hybrid journals, according to the accompanying ESAC Market Watch [^3] https://esac-initiative.org/market-watch/. Library consortia reported increased open access volume, streamlined payment and monitoring procedures, as well as extensive utilisation of open access options by the researchers they serve (Marques & Stone, 2020; Parmhed & Säll, 2023; Pinhasi et al., 2020). The ongoing standardisation of transformative agreements has contributed to improved transparency in terms of contracts and publisher-provided article metadata (Marques et al., 2019; Pinhasi et al., 2021). However, with the growing trend toward transformative agreements, continued reliance on big deals is perceived as problematic, because it perpetuates market concentration (Butler et al., 2023; Shu & Larivière, 2023). Whether transformative agreements lead to reduced pricing remains uncertain (Borrego, 2023) and a substantial transition of hybrid journals towards full open access could not be observed (Matthias et al., 2019; Momeni et al., 2021). The focus on large commercial publishers might also increase inequality (Ross-Hellauer et al., 2022), because transformative agreements’ focus on pay to publish mainly targets institutions from high-income countries, furthering a questionable journal prestige culture (Babini et al., 2022). Besides, an editorial-board resignation raised concerns that transformative agreements might encourage publishers to maximize journal publication volume “without regard to quality” (Rasmussen, 2023).
@@
The controversies surrounding hybrid open access and transformative agreements have led to varying policy conclusions. For instance, the Association of Swedish Higher Education Institutions (Sveriges universitets- och högskoleförbund, SUHF) recommended only supporting agreements for publishing in fully open access journals by 2026  [^4] https://www.su.se/english/news/open-access-need-to-move-away-from-transformative-agreements-1.683787. Likewise, most cOAlition S funders will discontinue financial support for transformative agreements by the end of 2024 (Liverpool, 2023). cOAlition S also removed the majority of hybrid journals from its Transformative Journal program in 2023 due to publishers’ failure to meet self-defined open access growth targets (Brainard, 2023). By contrast, the German DEAL consortium announced a five-year transformative agreement with Elsevier starting in 2024 and also renewed its contracts with Springer Nature and Wiley until the end of 2028. Similarly, the Colombia Consortium signed the first transformative agreements in Latin America (Muñoz-Vélez et al., 2024).
@@
Despite these controversies around transformative agreements as a means of transitioning journal publishing to full open access, there is limited evidence available on the uptake of open access in hybrid journals, and the extent to which this can be attributed to transformative agreements. Previous studies have focused on specific countries (Brayman et al., 2024; Haucap et al., 2021; Huang et al., 2020; Pölönen et al., 2020; Taubert et al., 2023; Wenaas, 2022) or publisher portfolios (Bakker et al., 2024; Fraser et al., 2023; Jahn et al., 2022; Momeni et al., 2023; Pieper & Broschinski, 2018; Schmal, 2024), while large-scale studies relied on self-reported agreement data (Moskovkin et al., 2022), or used APC pricing lists (Shu & Larivière, 2023). In particular, data availability is a limiting factor when studying the impact of transformative agreements (Bakker et al., 2024), because bibliometric databases, even though many allow the retrieval of open access articles in hybrid journals, do not directly attribute them to specific transformative agreements. Likewise, article-level open access invoicing and cost data are only partly available (Jahn et al., 2022; Kramer, 2024).
@@ 2 Methods
This study combines data from multiple publicly available sources, as shown in Figure 1. Initially, transformative agreement data retrieved from cOAlition S Journal Checker Tool [^5] https://www.coalition-s.org/blog/enabling-accurate-results-within-the-journal-checker-tool/ provided information about journal portfolios and participating institutions. After identifying hybrid journals by excluding fully open access journals, Crossref (Hendricks et al., 2020) served as the primary data source for article-level metadata including Creative Commons (CC) license information to indicate open access availability on publisher websites. To determine open access articles published through transformative agreements, first author affiliations from OpenAlex (Priem et al., 2022) were subsequently linked to eligible institutions according to the transformative agreement data. In the following, the steps are described in detail.
@@ 2.1 Transformative agreement and hybrid journal data
Data gathering started with obtaining journals included in transformative agreements from the publicly available Transformative Agreement Data dump [^6] https://journalcheckertool.org/transformative-agreements/ used by the cOAlition S Journal Checker Tool. [^7] https://www.coalition-s.org/blog/enabling-accurate-results-within-the-journal-checker-tool/ The dump consists of multiple online Google spreadsheets where each data file represents one agreement listed in the ESAC Transformative Agreement Registry. [^8] https://esac-initiative.org/about/transformative-agreements/agreement-registry/ From the retrieved spreadsheet files, journals and institutions involved per agreement were obtained.
A limitation of using the Journal Checker Tool and its underlying publicly available data dump to study the development of transformative agreements over time is that expired transformative agreements are constantly removed. To address this, four different snapshots were safeguarded and combined for this study: self-archived versions from July 2021, July 2022, and May 2023, as well as the most current dump downloaded on 11 December 2023. This ensured that transformative agreements, which ended from 2021 onwards, were included, representing the majority of transformative agreements. Overall, the four combined Transformative Agreement Data dumps used in this study contained 729 of 869 agreements listed in the ESAC Transformative Agreement Registry by December 2023.
@@ 2.2 Article and author metadata
Open access articles in hybrid journals were identified using the Creative Commons (CC) license information in Crossref metadata. License information relative to the “accepted manuscript (AM)” version was not considered. Crossref was used for open access identification because transformative agreement workflows generally require publishers to deliver CC license information to this DOI registration agency (Geschuhn & Stone, 2017).
@@
Affiliation metadata about corresponding authors are crucial for the planning and evaluation of transformative agreements because they are considered responsible for arranging open access publication (Borrego et al., 2021; Geschuhn & Stone, 2017; Schimmer et al., 2015). For this study, country and institutional affiliations were retrieved from OpenAlex. However, because corresponding author affiliation was only available for 54% of the investigated articles in OpenAlex, this study focused on first authors and their affiliations instead; approximately 90% of the investigated articles had first author affiliation metadata in OpenAlex, which is a much larger proportion than previously reported for the October 2022 snapshot (Zhang et al., 2024). First authors typically contribute the most to a paper and are often considered being lead authors (Larivière et al., 2016). Related studies also assumed first authors as a proxy to examine open access payments and transformative agreements (Haucap et al., 2021; Shu & Larivière, 2023; Zhang et al., 2022).
To estimate the impact of transformative agreements on hybrid open access, participating institutions from the Transformative Agreement Data dump, which the cOAlition S crowd-sourced from agreement documents and consortia, were matched with the first author affiliations recorded by OpenAlex using the ROR-ID. Matching also considered the duration of agreements according to the ESAC registry. Upon inspection, Transformative Agreement Data did not comprehensively cover associated institutions, such as university hospitals or institutes of large research organisations such as the Max Planck Society. To improve the matching, Transformative Agreement Data were automatically enriched with associated organisations using OpenAlex’s institution entity data.
The so compiled data set consists of 8,922,146 articles published in 12,857 hybrid journals between 2018 and 2022 (see Figure 2B). Hybrid journals in transformative agreements represented 40% of the total global output over the same time period, according to Crossref, while fully open access journals recorded 35%.
Figure 2: Initial data characteristics. (A) Full open access journals included in transformative agreements by evidence source Directory of Open Access Journals (DOAJ), OpenAlex and the Bielefeld GOLD OA list. (B) Number of articles in Crossref by journal types. The blue bars show the overall article volume of hybrid journals in transformative agreements, which were initially included in the study, in comparision with full open access journals according to OpenAlex. The remainder represents subscription-based journals not covered by transformative agreements.
@@ 3.1 Overview
Between 2018 and 2022, 11,189 out of 12,857 hybrid journals in transformative agreements published at least one open access article under a Creative Commons license. During this period, these hybrid journals provided open access to 742,369 out of 8,146,958 articles, representing a five-year open access proportion of 9.1%. Authors who could make use of transformative agreements at the time of publication contributed 328,957 open access articles. Overall, 394 transformative agreements recorded at least one open access article from an eligible institution.
@@
Figure 3A shows a moderate growth in the proportion of open access articles in hybrid journals, comparing the overall open access uptake and the impact of transformative agreements on this trend. Over the five-years period from 2018 to 2022, open access increased from 4.3% (n = 65,486) to 15% (n = 249,511). Simultaneously, the total article volume of the investigated journals rose from 1,528,051 in 2018 to 1,676,928 in 2022.
Figure 3B highlights that the majority of hybrid open access was made available through transformative agreements in 2021 and 2022. In 2022, 143,615 out of 249,511 open access articles were from eligible authors, representing 58%. However, there was also a notable increase in open access provision through other means, presumably publication fees being not invoiced through transformative agreements, from 4.1% (n = 62,625) in 2018 to 6.3% (n = 105,896) in 2022.
Figure 3C depicts the substantial variations among the hybrid journals included in transformative agreements in terms of open access uptake. Although the median generally follows the trend shown in Figure 3A, the farther stretch of upper quartiles and whiskers over the years illustrates that an increasing number of journals published an above-average proportion of open access articles. In 2022, 25% of hybrid journals (n = 2,576) had an open access uptake of 29%, and 6.6% of journals (n = 744) provided the majority of their articles under a Creative Commons license in the same year. On average, these journals were smaller (M = 75, SD = 186) than those with an open access share below 50% (M = 164, SD = 347).
@@ 3.2 Publishing market
Analysing hybrid open access across publishers between 2018 and 2022 reveals a large market concentration. Although 48 publishers offered transformative agreements, the three commercial publishers Elsevier, Springer Nature, and Wiley, accounted for 49% of the total article volume published (see Table 1). Together, they published 500,878 or 66% of the open access articles in hybrid journals. Elsevier, Springer Nature, and Wiley made 243,891 articles open access in hybrid journals through transformative agreements, resulting in an even larger market share of 74%.
@@
However, there are notable differences between the three large publishers. Although Elsevier published the largest volume of articles (n = 2,770,826, 34%), it recorded a comparably low number of open access articles, including those that are associated with transformative agreements. In contrast, Springer Nature and Wiley provided open access to a larger proportion of their articles (13% of Springer Nature articles and 15% of Wiley articles were open access), leading to higher open access market share (23% Springer Nature resp. 23% Wiley). This difference between Elsevier on the one hand and Springer Nature and Wiley on the other can be attributed to transformative agreements, as the latter made the majority of their open access articles available through such deals (57% Springer Nature resp. 55% Wiley).
@@
Figure 4 takes a closer look into the growth of hybrid open access across publishers by year, with a focus on open access enabled by transformative agreements. Although all publishers show a general long-term trend towards transformative agreements, Figure 4A and B indicate that Wiley experienced a substantial increase in its open access share from 5.9% (n = 11,628) in 2018 to 26% (n = 53,503) in 2022, representing a 4.5-fold increase. In contrast, Elsevier’s hybrid journals demonstrated a more modest increase, from 3.3% (n = 16,872) in 2018 to 10% (n = 60,821) in 2022, which is a relatively low open access share compared to the general trend. In 2018, Springer Nature had the largest open access proportion among the three publishers of 8.4% (n = 19,701), but experienced a relatively slower growth, resulting in 18% (n = 52,616) of articles being open access in Springer Nature hybrid journals in 2022.
@@
Figure 4: Developement of open access in hybrid journals included in transformative agreements between 2018 and 2022 by publishers. The blue areas represent open access through transformative agreements, the grey areas depict open access articles where no link to an agreement could be established. (A) Proportion of open access articles in hybrid journals per year and publisher. (B) Percentage of hybrid open access via agreements per year and publisher. Boxplots (C) show individual open access uptake rates by individual hybrid journals and open access funding per publication year and publisher. The individual outliers are not shown. Note that data on transformative agreements ending before June 2021 were not available for this study.
@@
The varying degrees of adoption of open access across the three major publishers can be attributed to distinct approaches to transformative agreements. Springer Nature, for example, started offering open access agreements for its hybrid journal portfolio to selected consortia such as the Max Planck Society, the Swedish Bibsam Consortium and the Finnish FinELib consortium in 2015 under the name Springer Compact [^10] https://web.archive.org/web/20180414062853id_/http://www.liber2015.org.uk/wp-content/uploads/2015/03/Springer-Compact.pdf. Prior to this, Springer had some pilot agreements with a small number of institutions, including the University of Göttingen (Schmidt & Shearer, 2012). However, the Springer Compact agreements were not included in the data as they concluded before the start of the transformative agreement data collection in June 2021. Nonetheless, the results demonstrate the importance of agreements for Springer Nature’s hybrid open access business over the past five years (Figure 2B). In 2022, 66% (n = 34,725) of open access in Springer Nature hybrid journals was enabled through transformative agreements. In the same year, 70% (n = 37,316) of Wiley’s open access articles could be linked to transformative agreements in 2022. By contrast, Elsevier published fewer than half of its open access articles through transformative agreements (n = 32,627; 54%).

### v3

# How open are hybrid journals included in transformative agreements?

@@ Abstract
The ongoing controversy surrounding transformative agreements, which aim to transition subscription-based journal publishing to full open access, highlights the need for large-scale studies assessing the impact of these agreements on hybrid open access. By combining multiple open data sources, including cOAlition S Journal Checker, Crossref, and OpenAlex, this study presents a novel approach that analyses over 700 agreements. Results suggest a strong growth in open access, from 4.3% in 2018 to 15% in 2022. Over five years, 11,189 hybrid journals provided open access to 742,369 out of 8,146,958 articles (9.1%). Authors who could make use of transformative agreements contributed 328,957 open access articles (44%) during this period, reaching a peak in 2022 with 143,615 out of 249,511 open access articles (58%). While this trend was predominantly driven by the three commercial publishers Elsevier, Springer Nature, and Wiley, open access uptake varied substantially across journals, publishers, disciplines, and countries. Particularly, the OECD and BRICS areas revealed different publication trends. In conclusion, this study suggests that current levels of implementation of transformative agreements is insufficient to bring about a large-scale transition to full open access.
@@ 1 Introduction
As a consequence, libraries and their consortia began to develop licensing strategies aimed at avoiding such ‘double dipping’ scenarios, in which well-established commercial publishers gain twice from reading and open access publishing fees (Mittermaier, 2015), as well as to increase publisher-provided immediate open access (Björk & Solomon, 2014; Schimmer et al., 2015). These considerations resulted in transformative agreements [^1] In this paper I use the term “transformative agreement”, addressing also offsetting, read-and-publish or publish-and-read deals, and other variants (Borrego et al., 2021; Hinchliffe, 2019). Although the term is critised as misleading and not useful to describe the different types of open access agreements between library consortia and commercial publishers (Babini et al., 2022), it is widely used in policy discussions and in the research literature., which cover a broad range of contracts between library consortia and publishers from the mid-2010s onwards, where institutional spending for subscriptions and open access publishing are considered together (Borrego et al., 2021; Hinchliffe, 2019). Transformative agreements seek to control costs while allowing a transitional phase for publishing more open access articles. Similar to big deals, transformative agreements mainly bundle hybrid and subscription-only journals from commercial publishers, but aim at a higher degree of transparency than previous big deals, where contracts including payments were confidential (Bergstrom et al., 2014).
@@
By the end of 2023, many transformative agreements had been implemented, but the interim outcomes were mixed. The ESAC Transformative Agreement Registry [^2] https://esac-initiative.org/about/transformative-agreements/agreement-registry/, the largest resource for library consortia to voluntarily and publicly share their agreements, recorded more than 800 transformative agreements. These agreements resulted in the publication of up to 900.000 open access articles published in both fully open access and hybrid journals, according to the accompanying ESAC Market Watch [^3] https://esac-initiative.org/market-watch/. Library consortia reported increased open access volume, streamlined payment and monitoring procedures, as well as extensive utilisation of open access options by the researchers they serve (Marques & Stone, 2020; Parmhed & Säll, 2023; Pinhasi et al., 2020). The ongoing standardisation of transformative agreements has contributed to improved transparency in terms of contracts and publisher-provided article metadata (Marques et al., 2019; Pinhasi et al., 2021). However, with the growing trend toward transformative agreements, continued reliance on big deals is perceived as problematic, because it perpetuates market concentration (Butler et al., 2023; Shu & Larivière, 2023). Whether transformative agreements lead to reduced pricing remains uncertain (Borrego, 2023) and a substantial transition of hybrid journals towards full open access could not be observed (Matthias et al., 2019; Momeni et al., 2021). The focus on large commercial publishers might also increase inequality (Ross-Hellauer et al., 2022), because transformative agreements’ focus on pay to publish mainly targets institutions from high-income countries, furthering a questionable journal prestige culture (Babini et al., 2022). Besides, an editorial-board resignation raised concerns that transformative agreements might encourage publishers to maximize journal publication volume “without regard to quality” (Rasmussen, 2023).
@@
The controversies surrounding hybrid open access and transformative agreements have led to varying policy conclusions. For instance, the British Joint Information Systems Committee (JISC) evaluated its open access strategy, which included transformative agreements (Brayman et al., 2024). The evaluation revealed that while these agreements had a significant impact on the growth of open access in the country, they had limited effects in facilitating a global shift towards full open access. As a result, the report advised that the British open access strategy should be reassessed. After making similar observation, Norwegian and Swedish universities and their consortia also argued for policy changes (Holden et al., 2023; Widding, 2024). Furthermore, cOAlition S conclude its financial support of transformative agreements at the end of 2024, but continue to view open access resulting from such agreements as compliant. [^4] https://www.coalition-s.org/coalition-s-confirms-the-end-of-its-financial-support-for-open-access-publishing-under-transformative-arrangements-after-2024/ cOAlition S also removed the majority of hybrid journals from its Transformative Journal program in 2023 due to publishers’ failure to meet self-defined open access growth targets (Brainard, 2023). By contrast, the German DEAL consortium announced a five-year transformative agreement with Elsevier starting in 2024 and also renewed its contracts with Springer Nature and Wiley until the end of 2028. Similarly, the Colombia Consortium signed the first transformative agreements in Latin America (Muñoz-Vélez et al., 2024).
@@
Despite these controversies around transformative agreements as a means of transitioning journal publishing to full open access, there is limited evidence available on the uptake of open access in hybrid journals, and the extent to which this can be attributed to transformative agreements. Previous studies have focused on specific countries (Brayman et al., 2024; Haucap et al., 2021; Huang et al., 2020; Pölönen et al., 2020; Taubert et al., 2023; Waltman & Lamers, 2022; Wenaas, 2022) or publisher portfolios (Bakker et al., 2024; Fraser et al., 2023; Jahn et al., 2022; Momeni et al., 2023; Pieper & Broschinski, 2018; Schmal, 2024), while large-scale studies relied on self-reported agreement data (Moskovkin et al., 2022), or used APC pricing lists (Shu & Larivière, 2023). In particular, data availability is a limiting factor when studying the impact of transformative agreements (Bakker et al., 2024), because bibliometric databases, even though many allow the retrieval of open access articles in hybrid journals, do not directly attribute them to specific transformative agreements. Likewise, article-level open access invoicing and cost data, which would make it possible to establish a direct link between transformative agreements and open access publications (Jahn et al., 2022; Kramer, 2024).
@@ 2 Methods
This study combines data from multiple publicly available sources, as shown in Figure 1. Initially, transformative agreement data retrieved from cOAlition S Journal Checker Tool [^5] https://www.coalition-s.org/blog/enabling-accurate-results-within-the-journal-checker-tool/ provided information about journal portfolios and participating institutions. After identifying hybrid journals by excluding fully open access journals, Crossref (Hendricks et al., 2020) served as the primary data source for article-level metadata including Creative Commons (CC) license information to indicate open access availability on publisher websites. Because of a lack of comprehensive publicly available invoicing data, open access articles published through transformative agreements were determined by linking first author affiliations from OpenAlex (Priem et al., 2022) to eligible institutions according to the transformative agreement data. In the following, the steps are described in detail.
@@ 2.1 Transformative agreement and hybrid journal data
Data gathering started with obtaining journals included in transformative agreements from the publicly available Transformative Agreement Data dump [^6] https://journalcheckertool.org/transformative-agreements/ used by the cOAlition S Journal Checker Tool [^7] https://www.coalition-s.org/blog/enabling-accurate-results-within-the-journal-checker-tool/, a voluntary effort based on publicly disclosed contracts in the ESAC Transformative Agreement Registry [^8] https://esac-initiative.org/about/transformative-agreements/agreement-registry/. The dump consists of multiple online Google spreadsheets where each data file represents one agreement listed in the ESAC Transformative Agreement Registry. From the retrieved spreadsheet files, journals and institutions involved per agreement were obtained.
It should be noted that although many library consortia see the need to register their agreements through the ESAC registry, some fail to disclose all details, including the full-text of contracts (Bakker et al., 2024; Kramer, 2024). Furthermore, the extent to which ESAC is comprehensive remains uncertain, potentially limiting the coverage of transformative agreements in the Journal Checker Tool. Another limitation of using the Journal Checker Tool and its underlying publicly available data dump to study the development of transformative agreements over time is that expired transformative agreements are constantly removed. To address this, four different snapshots were safeguarded and combined for this study: self-archived versions from July 2021, July 2022, and May 2023, as well as the most current dump downloaded on 11 December 2023. This ensured that transformative agreements, which ended from 2021 onwards, were included, representing the majority of transformative agreements. Overall, the four combined Transformative Agreement Data dumps used in this study contained 729 of 869 agreements listed in the ESAC Transformative Agreement Registry by December 2023.
@@ 2.2 Article and author metadata
Open access articles in hybrid journals were identified using the Creative Commons (CC) license information in Crossref metadata, with consideration of any CC variant. License information relative to the “accepted manuscript (AM)” version was not considered. Crossref was used for open access identification because transformative agreement workflows generally require publishers to deliver CC license information to this DOI registration agency (Geschuhn & Stone, 2017).
@@
Affiliation metadata about corresponding authors are crucial for the planning and evaluation of transformative agreements because they are considered responsible for arranging open access publication (Borrego et al., 2021; Geschuhn & Stone, 2017; Schimmer et al., 2015). For this study, country and institutional affiliations were retrieved from OpenAlex. Because the corresponding authorship field was not fully supported by OpenAlex at the time of analysis, and respective affiliation data were only available for 54% of the investigated articles, this study focused on first authors and their affiliations instead; approximately 90% of the articles examined had first author affiliation metadata in OpenAlex, which is a much larger proportion than previously reported for the October 2022 snapshot (Zhang et al., 2024). First authors typically contribute the most to a paper and are often considered as lead authors (Chinchilla-Rodríguez et al., 2024; Larivière et al., 2016). In case of lacking data about corresponding authors in bibliometric databases, related studies also utilised first authors as a proxy to examine open access payments and transformative agreements (Haucap et al., 2021; Shu & Larivière, 2023; Zhang et al., 2022).
To estimate the impact of transformative agreements on hybrid open access, participating institutions from the Transformative Agreement Data dump, which the cOAlition S crowd-sourced from agreement documents and consortia, were matched with the first author affiliations recorded by OpenAlex using the ROR-ID. Matching also considered the duration of agreements according to the ESAC registry. In total, 502 agreements were active between 2018 and 2022. Upon inspection, Transformative Agreement Data did not comprehensively cover associated institutions, such as university hospitals or institutes of large research organisations such as the Max Planck Society. To improve the matching, Transformative Agreement Data were automatically enriched with ROR-IDs from associated organisations according to OpenAlex’s institution entity data.
Figure 2: Initial data characteristics. (A) Full open access journals included in transformative agreements by evidence source Directory of Open Access Journals (DOAJ), OpenAlex and the Bielefeld GOLD OA list. (B) Number of articles in Crossref by journal types. The purple bars show the overall article volume of hybrid journals in transformative agreements, which were initially included in the study, in comparision with full open access journals according to OpenAlex. The remainder represents articles in subscription-based journals not covered by transformative agreements.
The so compiled data set consists of 8,922,146 articles published in 12,857 hybrid journals included in at least one transformative agreement between 2018 and 2022 (see purple bars in Figure 2B). These hybrid journals in transformative agreements represented 40% of the total global output over the same time period, according to Crossref, while fully open access journals recorded 35%.
@@ 3.1 Overview
Between 2018 and 2022, 11,189 out of 12,857 hybrid journals in transformative agreements published at least one open access article under a Creative Commons license. These eligible 11,189 hybrid journals constituted the foundation for the subsequent analyses. Collectively, they provided open access to 742,369 out of 8,146,958 articles during the investigated period, representing a five-year open access proportion of 9.1%. Authors who could make use of transformative agreements at the time of publication contributed 328,957 (44%) open access articles. Overall, this investigation was able to establish a link between open access articles and eligible institutions for 394 out of 502 (78%) active transformative agreements between 2018 and 2022 through the linking of first author affiliations with agreement data.
@@
Figure 3A shows a moderate growth in the proportion of open access articles in hybrid journals, comparing the overall open access uptake and the impact of transformative agreements on this trend. Over the five-years period from 2018 to 2022, open access in hybrid journals increased from 4.3% (n = 65,486) to 15% (n = 249,511). Simultaneously, the total article volume of the investigated journals rose from 1,528,051 in 2018 to 1,676,928 in 2022.
Figure 3B highlights that the majority of hybrid open access was made available through transformative agreements in 2021 and 2022. In 2022, 143,615 out of 249,511 open access articles were from eligible authors, representing 58%. However, there was also a notable increase in open access provision through other means, presumably publication fees being not invoiced through transformative agreements as indicated by the grey area in 3A, from 4.1% (n = 62,625) in 2018 to 6.3% (n = 105,896) in 2022.
Figure 3C depicts the substantial variations among the hybrid journals included in transformative agreements in terms of open access uptake. Although the median generally follows the trend shown in Figure 3A, the farther stretch of upper quartiles and whiskers over the years illustrates that an increasing number of journals published an above-average proportion of open access articles. In 2022, 25% of hybrid journals (n = 2,576) had an open access uptake of 29%, and 6.6% of journals (n = 744) provided the majority of their articles under a CC license in the same year. On average, these journals were smaller (M = 75, SD = 186) than those with an open access share below 50% (M = 164, SD = 347).
@@ 3.2 Publishing market
Analysing hybrid open access across publishers between 2018 and 2022 reveals a large market concentration. Although 48 publishers offered transformative agreements, the three commercial publishers Elsevier, Springer Nature, and Wiley accounted for 49% of hybrid journals, representing 5,144,308 or 63% of the total article volume (see Table 1). Together, they published 500,878 or 66% of the open access articles in hybrid journals. Elsevier, Springer Nature, and Wiley made 243,891 articles open access in hybrid journals through transformative agreements, resulting in an even larger market share of 74%.
@@
However, there are notable differences between the three large publishers. Although Elsevier published the largest volume of articles (n = 2,770,826; 34%), it recorded a comparably low number of open access articles, including those that are associated with transformative agreements. In contrast, Springer Nature and Wiley provided open access to a larger proportion of their articles (13% of Springer Nature articles and 15% of Wiley articles were open access), leading to higher open access market shares (23% Springer Nature resp. 23% Wiley). This difference between Elsevier on the one hand and Springer Nature and Wiley on the other can be attributed to transformative agreements, as the latter made the majority of their open access articles available through such deals (57% Springer Nature resp. 55% Wiley).
@@
Figure 4 takes a closer look into the growth of hybrid open access across publishers by year, with a focus on open access enabled by transformative agreements. Although all publishers show a general long-term trend towards transformative agreements, Figure 4A and B indicate that Wiley experienced a substantial increase in its open access share from 5.9% (n = 11,628) in 2018 to 26% (n = 53,503) in 2022. In contrast, Elsevier’s hybrid journals demonstrated a more modest increase, from 3.3% (n = 16,872) in 2018 to 10% (n = 60,821) in 2022, which is a relatively low open access share compared to the general trend. In 2018, Springer Nature had the largest open access proportion among the three publishers of 8.4% (n = 19,701), but experienced a relatively slower growth, resulting in 18% (n = 52,616) of articles being open access in Springer Nature hybrid journals in 2022.
@@
Figure 4: Development of open access in hybrid journals included in transformative agreements between 2018 and 2022 by publishers. The blue areas represent open access through transformative agreements, the grey areas depict open access articles where no link to an agreement could be established. (A) Proportion of open access articles in hybrid journals per year and publisher. (B) Percentage of hybrid open access via agreements per year and publisher. Boxplots (C) show individual open access uptake rates by individual hybrid journals and open access funding per publication year and publisher. The individual outliers are not shown. Note that data on transformative agreements ending before June 2021 were not available for this study.
@@
The varying degrees of adoption of open access across the three major publishers can be attributed to distinct approaches to transformative agreements. Springer Nature, for example, started offering open access agreements for its hybrid journal portfolio to selected consortia such as the Max Planck Society, the Swedish Bibsam consortium and the Finnish FinELib consortium in 2015 under the name Springer Compact [^10] https://web.archive.org/web/20180414062853id_/http://www.liber2015.org.uk/wp-content/uploads/2015/03/Springer-Compact.pdf. Prior to this, Springer had some pilot agreements with a small number of institutions, including the University of Göttingen (Schmidt & Shearer, 2012). However, the Springer Compact agreements were not included in the data as they concluded before the start of the transformative agreement data collection in June 2021. Nonetheless, the results demonstrate the importance of agreements for Springer Nature’s hybrid open access business over the five-years period (Figure 4B). In 2022, 66% (n = 34,725) of open access in Springer Nature hybrid journals was enabled through transformative agreements. In the same year, 70% (n = 37,316) of Wiley’s open access articles could be linked to transformative agreements. By contrast, Elsevier published a comparatively lower proportion of its open access articles through transformative agreements in 2022 (n = 32,627; 54%).